# A Simple Lifelong Learning Approach

### v19

@@ Abstract
In lifelong learning, data are used to improve performance not only on the present task, but also on past and future (unencountered) tasks. While typical transfer learning algorithms can improve performance on future tasks, their performance on prior tasks degrades upon learning new tasks (called forgetting). Many recent approaches for continual or lifelong learning have attempted to maintain performance on old tasks given new tasks. But striving to avoid forgetting sets the goal unnecessarily low. The goal of lifelong learning should be to use data to improve performance on both future tasks (forward transfer) and past tasks (backward transfer). In this paper, we show that a simple approach—representation ensembling—demonstrates both forward and backward transfer in a variety of simulated and benchmark data scenarios, including tabular, vision (CIFAR-100, 5-dataset, Split Mini-Imagenet, and Food1k), and speech (spoken digit), in contrast to various reference algorithms, which typically failed to transfer either forward or backward, or both. Moreover, our proposed approach can flexibly operate with or without a computational budget.
1Johns Hopkins University, 2Baylor College of Medicine, 3University of Cambridge, 4Microsoft Research, denotes equal contribution, corresponding author: jdey4@jhu.edu
1 Introduction
@@
Learning is a process by which an intelligent system improves performance on a given task by leveraging data [1]. In classical machine learning, the system is often optimized for a single task [2, 3]. While it is relatively easy to simultaneously optimize for multiple tasks (multi-task learning) [4], it has proven much more difficult to sequentially optimize for multiple tasks [5, 6]. Specifically, classical machine learning systems, and natural extensions thereof, exhibit “catastrophic forgetting” when trained sequentially, meaning their performance on the prior tasks drops precipitously upon training on new tasks [7, 8]. However, learning could be lifelong, with agents continually building on past knowledge and experiences, improving on many tasks given data associated with any task. For example, in humans, learning a second language often improves performance in an individual’s native language [9].
In the past 30 years, a number of sequential task learning algorithms have attempted to overcome catastrophic forgetting. These approaches naturally fall into one of two camps. In one camp, the algorithm has fixed resources, and so must reallocate resources (essentially compressing representations) in order to incorporate new knowledge [10, 11, 12, 13, 14]. Biologically, this corresponds to adulthood, where brains have a nearly fixed or decreasing number of cells and synapses. In the other camp, the algorithm adds (or builds) resources as new data arrive (essentially ensembling representations) [15, 16, 17, 18]. Biologically, this corresponds to development, where brains grow by adding cells, synapses, etc.
Approaches from both camps demonstrate some degree of continual (or lifelong) learning [19]. In particular, they can sometimes learn new tasks faster due to prior learning on related tasks, while not catastrophically forgetting old tasks (see Appendix .1 for a detailed discussion on the relevant algorithms). However, as we will show, many lifelong learning algorithms are unable to transfer knowledge forward (to future unseen tasks) and most of them do not transfer backward (to previously seen tasks). With high enough sample sizes, some of these algorithms are able to transfer forward or backward, but transfer is more important in low sample size regimes [20, 17]. This inability to effectively transfer in low-sample size regimes has been identified as one of the key obstacles limiting the capabilities of artificial intelligence [21, 22].
@@
In this paper, we propose a general approach for lifelong learning which can be used with many existing encoder models. Specifically, we focus our approach on ensembling deep networks (Simple Lifelong Learning Networks, SiLLy-N). Additionally, we demonstrate how the same approach can be generalized for lifelong learning based on ensembling decision forests (Simple Lifelong Learning Forests, SiLLy-F). We explore our proposed algorithm as compared to a number of reference algorithms on an extensive suite of numerical experiments that span simulation, vision datasets including CIFAR-100, 5-dataset, Split Mini-Imagenet, and Food1k, as well as the spoken digit dataset. Figure 1 illustrates that our algorithm outperforms all the reference algorithms in terms of forward, backward, and overall transfer on CIFAR 10X10 dataset. Ablation studies indicate the degree to which the amount of representation or storage capacity and replaying old task data impact performance of our algorithms. All our code and experiments are open source to facilitate reproducibility.
@@
Figure 1: Performance summary on CIFAR 10X10 benchmark dataset. Columns are different evaluation criteria (see Section 2 for definitions, and Section 6 for experimental details), each strip of colored dots corresponds to an algorithm (we introduce SiLLy-N here) and each dot represents a task. In all the figure, older tasks have darker colors. Resource growing algorithms have a ‘*’. EWC, O-EWC, SI, TAG and ER always perform worse than LwF, and hence we do not show them in the plot. SiLLy-N (red) outperforms all reference algorithms in terms of forward (second panel), backward (third panel), and overall transfer (first panel). Importantly, such better transfer is achieved at high overall accuracy (last panel). More datasets are evaluated in Figure 8.
2 Mathematical Framework
2.1 The lifelong learning objective
@@
Consider a lifelong or continual learning environment with tasks, $\mathcal{T}=\{1,2,\cdots,T\}$. We consider task aware lifelong learning, i.e., the tasks are known during both training and testing time. For simplicity, we consider that any task $t\in\mathcal{T}$ has the same input space, i.e., they have $\mathcal{X}\subset\mathbb{R}^{D}$ valued inputs with $\mathcal{Y}=\{1,\cdots,K_{t}\}$ valued class labels. We assume the tasks arrive sequentially, but the $n_{t}$ data samples within each task $t$, $\mathbf{S}^{t}=\{(X_{i},Y_{i})\}_{i=1}^{n_{t}}$, are batched and sampled identically and independently ($iid$) from some fixed distribution $\mathcal{D}_{t}$, and $\sum_{t=1}^{T}n_{t}=n$. A learner $f\in\mathcal{F}$ trains on $\mathbf{S}^{t}$ and chooses a hypothesis $h_{t}\in\mathcal{H}$ associated with task $t$ by minimizing a particular risk, where $\mathcal{F}$ and $\mathcal{H}$ are the algorithm and the hypothesis space, respectively. In supervised learning settings, one can consider the following risk for a particular task $t$:
(1) $R^{t}(f(\mathbf{S}^{t}))=R^{t}(h_{t})=\mathbb{E}_{(X,Y)\sim\mathcal{D}_{t}}[\ell_{t}(h_{t}(X),Y)],$
where $\ell_{t}:\mathcal{Y}\times\mathcal{Y}\rightarrow[0,\infty)$ is a given loss function associated with the task $t$ and $h_{t}=f(\mathbf{S}^{t})$. Note that the data $\mathbf{S}^{t}$ may contain data that is relevant to any number of tasks (potentially all the tasks) in the environment. One may take expectation over $\mathbf{S}^{t}$ and consider the generalization error for the task as:
(2) $\mathcal{E}^{t}_{f}(\mathbf{S}^{t})=\mathbb{E}[R^{t}(f(\mathbf{S}^{t}))].$
In the above equation, the learner will have access to a total of $T$ datasets after $T$ tasks, $\bigcup_{t=1}^{T}\mathbf{S}^{t}$, instead of $\mathbf{S}^{t}$ only [^1] More generally, we may have $J$ datasets, where $J\neq T$ and each dataset may be associated with the target distributions of multiple tasks. For simplicity, we do not consider such scenarios further at this time.. The goal is to find a learner $f\in\mathcal{F}$ that chooses a set of total $T$ hypotheses $\{h_{1},\cdots,h_{T}\}$ (one hypothesis for each task) such that the generalization error over all the tasks after observing all the data is minimized, that is:
@@
(3) $\begin{array}[]{ll}\textnormal{minimize}&\sum_{t=1}^{T}\mathcal{E}_{f}^{t}(\bigcup_{t^{\prime}=1}^{T}\mathbf{S}^{t^{\prime}})\\ \textnormal{subject to }&f\in\mathcal{F}\end{array}\enspace.$
2.2 Lifelong learning evaluation criteria
@@
(4) $\mathsf{Transfer}^{t}(f):=\log\frac{\mathcal{E}^{t}_{f}(\mathbf{S}^{t})}{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{T}\mathbf{S}^{t^{\prime}})}.$
We say that an algorithm $f$ has transferred to task $t$ from all the tasks up to $T$ if and only if $\mathsf{Transfer}^{t}(f)>0$.
@@
Definition 2 (Forward Transfer).
@@
(5) $\mathsf{Forward~{}Transfer}^{t}(f):=\log\frac{\mathcal{E}^{t}_{f}(\mathbf{S}^{t})}{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{t}\mathbf{S}^{t^{\prime}})}.$
We say an algorithm (positively) forward transfers for task $t$ if and only if:
(6) $\mathsf{Forward~{}Transfer}^{t}(f)>0.$
@@
Definition 3 (Backward Transfer).
@@
(7) $\mathsf{Backward~{}Transfer}^{t}(f):=\log\frac{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{t}\mathbf{S}^{t^{\prime}})}{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{T}\mathbf{S}^{t^{\prime}})}.$
We say an algorithm backward transfers to Task $t$ from all the future tasks up to $T$ if and only if:
(8) $\mathsf{Backward~{}Transfer}^{t}(f)>0.$
Note that $\mathsf{Transfer}$ can be decomposed into $\mathsf{Forward~{}Transfer}$ and $\mathsf{Backward~{}Transfer}$:
(9) $\mathsf{Transfer}^{t}(f)$ $=\log\frac{\mathcal{E}^{t}_{f}(\mathbf{S}^{t})}{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{T}\mathbf{S}^{t^{\prime}})}$
(10) $=\log\frac{\mathcal{E}^{t}_{f}(\mathbf{S}^{t})}{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{t}\mathbf{S}^{t^{\prime}})}+\log\frac{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{t}\mathbf{S}^{t^{\prime}})}{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{T}\mathbf{S}^{t^{\prime}})}$
(11) $=$ $\mathsf{Forward~{}Transfer}^{t}(f)+\mathsf{Backward~{}Transfer}^{t}(f).$
Another paper [26], concomitantly introduced transfer and forgetting (backward transfer). Their statistics are the same as ours, except they do not use a $\log$. We opted for a $\log$ to address numerical stability issues in comparing small numbers. Because log is a monotonic function, the order of ranking algorithms is preserved (Appendix Figure 1 shows a version of Figure 1 and Figure 8, but using Veniat’s statistics, which is nearly visually identical). By virtue of introducing $\mathsf{Forward~{}Transfer}$ here, we can identify the inherent trade-off between forward and backward transfer, for a fixed amount of total transfer. Apart from the above statistics, we also report accuracy per task.
Definition 4 (Accuracy).
@@
(12) $\mathsf{Accuracy}^{t}(f):=1-{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{T}\mathbf{S}^{t^{\prime}})}.$
3 Representation ensembling algorithms

### v21

# Simple Lifelong Learning Machines

@@ Abstract
In lifelong learning, data are used to improve performance not only on the present task, but also on past and future (unencountered) tasks. While typical transfer learning algorithms can improve performance on future tasks, their performance on prior tasks degrades upon learning new tasks (called forgetting). Many recent approaches for continual or lifelong learning have attempted to maintain performance on old tasks given new tasks. But striving to avoid forgetting sets the goal unnecessarily low. The goal of lifelong learning should be to use data to improve performance on both future tasks (forward transfer) and past tasks (backward transfer). In this paper, we show that a simple approach—representation ensembling—demonstrates both forward and backward transfer in a variety of simulated and benchmark data scenarios, including tabular, vision (CIFAR-100, 5-dataset, Split Mini-Imagenet, Food1k, and CORe50), and speech (spoken digit), in contrast to various reference algorithms, which typically failed to transfer either forward or backward, or both. Moreover, our proposed approach can flexibly operate with or without a computational budget.
Index Terms:
continual/lifelong learning, forward transfer, backward transfer, replay
I Introduction
@@
Learning is a process by which an intelligent system improves performance on a given task by leveraging data [1]. In classical machine learning, the system is often optimized for a single task [2, 3]. While it is relatively easy to simultaneously optimize for multiple tasks (multi-task learning) [4], it has proven much more difficult to sequentially optimize for multiple tasks [5, 6]. Specifically, classical machine learning systems, and natural extensions thereof, exhibit “catastrophic forgetting” when trained sequentially, meaning their performance on the prior tasks drops precipitously upon training on new tasks [7, 8, 9]. However, learning could be lifelong, with agents continually building on past knowledge and experiences, improving on many tasks given data associated with any task. For example, in humans, learning a second language often improves performance in an individual’s native language [10].
In lifelong learning, where tasks arrive sequentially, the ability to transfer knowledge across tasks is characterized by two complementary objectives: forward transfer and backward transfer. Forward transfer facilitates accelerated learning in new tasks using previous knowledge. In contrast, backward transfer evaluates the impact of new learning on previously encountered tasks. Achieving both positive forward and backward transfer is crucial for an effective lifelong learner. However, as we will demonstrate, many existing lifelong learning algorithms do not enable forward transfer to future tasks, and most do not exhibit positive backward transfer to previously learned tasks.
@@
In this paper, we propose a general and simple approach for lifelong learning which can be used with many existing encoder models. Specifically, we focus our approach on ensembling deep networks (Simple Lifelong Learning Networks, SiLLy-N). Additionally, we demonstrate how the same approach can be generalized for lifelong learning based on ensembling decision forests (Simple Lifelong Learning Forests, SiLLy-F). Table I and Figure 3 show our proposed approaches grow linearly with the task number and can flexibly operate with both growing and constant resources depending on the available computation budget. Moreover, we explore our proposed algorithm as compared to a number of reference algorithms on an extensive suite of numerical experiments that span simulation, vision datasets including CIFAR-100, 5-dataset, Split Mini-Imagenet, and Food1k, as well as the spoken digit dataset. Figure 1 illustrates that our algorithm outperforms all the reference algorithms in terms of forward, backward, and overall transfer on different vision and speech datasets. Ablation studies indicate the degree to which the amount of representation or storage capacity and replaying old task data impact performance of our algorithms. All our code and experiments are open source to facilitate reproducibility.
The subsequent organization of the paper can be summarized as:
  1. 1.
    Section II presents a discussion of relevant algorithms, highlighting their architecture and computational complexity in comparison to SiLLy-N and SiLLy-F.
  2. 2.
    Section III introduces the learning environment and proposes three transfer statistics.
  3. 3.
    Section IV details the design of SiLLy-N and SiLLy-F, incorporating insights from various ensembling approaches.
  4. 4.
    Sections V and VI provide empirical evaluations of SiLLy-N and SiLLy-F on both simulated and benchmark datasets.
  5. 5.
    Section VII concludes with a discussion of strengths, limitations, and directions for future research.
@@
Figure 1: Performance summary on different vision and speech benchmark datasets. Columns are different evaluation criteria (see Section III for definitions, and Section VI for experimental details), each strip of colored dots corresponds to an algorithm (we introduce SiLLy-N here) and each dot represents a task. Older tasks have darker colors. Resource growing algorithms have a ‘*’. EWC, O-EWC, SI, TAG and ER always perform worse than LwF, and hence we do not show them in the plot. SiLLy-N (red) outperforms all reference algorithms in terms of overall (first column), forward (second column), and backward (third column). Importantly, such better transfer is achieved at high overall accuracy (fourth column). For CORe50 dataset, we are unable to run Model Zoo in the resource constrained environment of the experiment and LwF completely forgets the first several tasks with the task-accuracies going down to the chance level ($20\%$). See Appendix Figure 1 for the error bars for each dot in the figure.
II Related Works
In this work, we propose a lifelong learning approach based on representation ensembling which can flexibly operate in both resource growing and resource constrained modes. Existing resource constrained algorithms, such as EWC [11], Online EWC [12], EWC++ [13], SI [14], and LwF [15], use regularization techniques to exploit the stability-plasticity trade-off to mitigate forgetting.
On the contrary, the resource growing approaches add resources as they face more tasks. Authors from [16] used a weighted ensemble of learners in a streaming setting with distribution shift. TrAdaBoost [17] boosts ensemble of learners to enable transfer learning. In continual learning scenarios, many algorithms have been built on these ideas by ensembling dependent representations. For example, Learn++ [18] boosts ensembles of weak learners learned over different data sequences in class incremental lifelong learning settings [19]. Model Zoo [20] uses the same boosting approach in task incremental lifelong learning scenarios. Another group of algorithms, ProgNN [21] and DF-CNN [22] learn a new “column” of nodes and edges with each new task, and ensembles the columns for inference (such approaches are commonly called ‘modular’ now). The primary difference between ProgNN and DF-CNN is that ProgNN has forward connections to the current column from all the past columns. This creates the possibility of forward transfer while freezing backward transfer. However, the forward connections in ProgNN render it computationally inefficient for a large number of tasks. DF-CNN gets around this problem by learning a common knowledge base and thereby, creating the possibility of backward transfer.
Recently, many other modular approaches have been proposed in the literature that improve on ProgNN’s capacity growth. These methods consider the capacity for each task being composed of modules that can be shared across tasks and grown as necessary. For example, PackNet [23] starts with a fixed capacity network and trains for additional tasks by freeing up portion of the network capacity using iterative pruning. Veniat et al. [24] trains additional modules with each new task, and the old modules are only used selectively. [25] improved the memory efficiency of the modular methods by adding new modules according to the complexity of the new tasks. Authors in [26] proposed non-parametric factorization of the layer weights that promotes sharing of the weights between tasks. However, all of modular methods described above lack backward transfer because the old modules are not updated with the new tasks. Dynamically Expandable Representation (DER) [27] proposed an improvement over the modular approaches where the model capacity is dynamically expanded and the model is fine-tuned by replaying a portion of the old task data along with the new task data. This approach achieves backward transfer between tasks as reported by the authors in the experiments. Another modular approach, proposed by Kang et al. [28], uses sub-networks inspired by the lottery ticket hypothesis [29] and dynamically expands the network capacity with the number of tasks.
Another strategy for building lifelong learning machines is to use total or partial replay [30, 31, 32, 33, 34]. Replay approaches keep the old data and replay them when faced with new tasks to mitigate catastrophic forgetting. Many recent approaches combine replay with other transfer learning strategies like knowledge distillation [35] to improve storage efficiency of replay data. Recently, hypernetwork based old network weight generation accompanied by synthetic replay [36] demonstrated potential of memory-efficient operation in a learning environment with a huge number of tasks. Another replay algorithm, MER [37], aligns the gradient update between task datasets while replaying old task data. However, as we illustrated in the main text, previously proposed replay algorithms do not demonstrate positive backward transfer in our experiments, though they often do not forget as much as other approaches.
Our approach builds directly on previously proposed modular and replay approaches with one key distinction: in our approach, representations are learned independently. Independent representations also have computational advantages, as doing so merely requires quasilinear time and space, and can be learned in parallel.
III Mathematical Framework
III-A The lifelong learning objective
@@
In a lifelong learning setting, we consider a sequence of tasks $\mathcal{T}=\{1,2,\dots,T\}$, where the tasks are known during training and testing. Each task $t\in\mathcal{T}$ shares the same input space, $\mathcal{X}\subset\mathbb{R}^{D}$, with class labels $\mathcal{Y}=\{1,\dots,K_{t}\}$. The tasks arrive sequentially, and within each task $t$, the dataset $\mathbf{S}^{t}=\{(X_{i},Y_{i})\}_{i=1}^{n_{t}}$ is sampled $iid$ from a fixed distribution $\mathcal{D}_{t}$, with $\sum_{t=1}^{T}n_{t}=n$.
Given access to all observed datasets $\bigcup_{t=1}^{T}\mathbf{S}^{t}$, the goal is to find a learner $f$ that minimizes the overall generalization error across all tasks:
@@
$\begin{array}[]{ll}\textnormal{minimize}&\sum_{t=1}^{T}\mathcal{E}_{f}^{t}(\bigcup_{t^{\prime}=1}^{T}\mathbf{S}^{t^{\prime}})\\ \textnormal{subject to }&f\in\mathcal{F}\end{array}\enspace,$ (1)
where $\mathcal{E}_{f}^{t}$ is the generalization error or expected risk on task $t$ and the learner will have access to a total of $T$ datasets after $T$ tasks, $\bigcup_{t=1}^{T}\mathbf{S}^{t}$  [^1] More generally, we may have $J$ datasets, where $J\neq T$ and each dataset may be associated with the target distributions of multiple tasks. For simplicity, we do not consider such scenarios further at this time.. Risk is defined as the expected task-specific loss $\ell_{t}:\mathcal{Y}\times\mathcal{Y}\rightarrow[0,\infty)$ in task $t$ (see [38] for a detailed formulation on different out-of-distribution learning scenarios).
III-B Lifelong learning evaluation criteria
@@
$\mathsf{Transfer}^{t}(f):=\log\frac{\mathcal{E}^{t}_{f}(\mathbf{S}^{t})}{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{T}\mathbf{S}^{t^{\prime}})}.$ (2)
In words, Equation 2 quantifies the extent that a learner $f$, is able to improve the performance on task $t$, when using the data on all other tasks, $(1,\dots,T)$. It does so by computing the ratio of generalization error on task $t$ while only training on data from task $t$, relative to the generalization error on task $t$ when training on the data from all tasks.
@@
Definition 2 (Forward Transfer)
@@
$\mathsf{Forward~{}Transfer}^{t}(f):=\log\frac{\mathcal{E}^{t}_{f}(\mathbf{S}^{t})}{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{t}\mathbf{S}^{t^{\prime}})}.$ (3)
Forward transfer is identical to transfer as defined in Equation 2, except that it only considers the relation between training on data from task $t$ as compared to training on the prior tasks, $(1,\dots,t)$, excluding future tasks, $(t+1,\dots,T)$.
@@
Definition 3 (Backward Transfer)
@@
$\mathsf{Backward~{}Transfer}^{t}(f):=\log\frac{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{t}\mathbf{S}^{t^{\prime}})}{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{T}\mathbf{S}^{t^{\prime}})}.$ (4)
Backward transfer is also like Equation 2, except it compares the performance when training on all prior tasks, $(1,\dots,t)$, with training on all tasks, $(1,\dots,T)$.
Note that $\mathsf{Transfer}$ can be decomposed into $\mathsf{Forward~{}Transfer}$ and $\mathsf{Backward~{}Transfer}$ (see Appendix -A for details):
$\mathsf{Transfer}^{t}(f)=\mathsf{Forward~{}Transfer}^{t}(f)+\mathsf{Backward~{}Transfer}^{t}(f).$ (5)
The above equation underscores the sequential progression of tasks in lifelong learning, indicating that the total knowledge a learner acquires is derived from both previously encountered tasks (forward transfer) and those yet to be encountered (backward transfer). While forward transfer is relatively straightforward to achieve [44], backward transfer presents a significant challenge due to the issue of catastrophic forgetting. In particular, a learner that demonstrates strong forward transfer ($\mathsf{Forward~{}Transfer}^{t}(f)>0$) but experiences negative backward transfer ($\mathsf{Backward~{}Transfer}^{t}(f)<0$) from all future tasks will eventually lose the learned knowledge, resulting in $\mathsf{Transfer}^{t}(f)\leq 0$ for the task.
Another paper [24], concomitantly introduced transfer and forgetting (backward transfer). Their statistics are the same as ours, except they do not use a $\log$. We opted for a $\log$ to address numerical stability issues in comparing small numbers. Because log is a monotonic function, the order of ranking algorithms is preserved (Appendix Figure 2 shows a version of Figure 1, but using Veniat’s statistics, which is nearly visually identical). By virtue of introducing $\mathsf{Forward~{}Transfer}$ here, we can identify the inherent trade-off between forward and backward transfer, for a fixed amount of total transfer. Apart from the above statistics, we also report accuracy per task.
Definition 4 (Accuracy)
@@
$\mathsf{Accuracy}^{t}(f):=1-{\mathcal{E}^{t}_{f}(\bigcup_{t^{\prime}=1}^{T}\mathbf{S}^{t^{\prime}})}.$ (6)
III-C Computational Taxonomy of Lifelong Learners
TABLE I: A computational taxonomy of lifelong learners. We show soft-O notation ($\tilde{\mathcal{O}}(\cdot,\cdot)$) as a function of total training samples, $n=\sum_{t}^{T}n_{t}$, where $n_{t}$ is the number of training samples for the $t^{th}$ task and total task, $T$, as well as the common setting where $n$ is proportional to $T$. Parametric, semi-parametric, and non-parametric algorithms have parameters that remain fixed, grow slowly, and scale proportionally to $n$, respectively.
Parametric Capacity Space Time Examples
($n,T$) ($n,T$) ($n\propto T$) $(n,T)$ $(n\propto T)$
parametric $1$ $1$ $1$ $1$ $1$ SiLLy-N-M, SiLLy-F-M
parametric $1$ $1$ $1$ $n$ $n$ O-EWC [12], SI [14], LwF [15]
parametric $1$ $T$ $n$ $nT$ $n^{2}$ EWC [11]
parametric $1$ $n$ $n$ $nT$ $n^{2}$ Total Replay
semi-parametric $T$ $T^{2}$ $n^{2}$ $nT$ $n^{2}$ ProgNN[21]
semi-parametric $T$ $T$ $n$ $n$ $n$ DF-CNN [22]
semi-parametric $T$ $T+n$ $n$ $n$ $n$ SiLLy-N, Model Zoo[20], DER [27], LMC[25]
non-parametric $n$ $n$ $n$ $n$ $n$ SiLLy-F, IBP-WF [26]
In the past 30 years, a number of lifelong learning algorithms have attempted to overcome catastrophic forgetting [45, 46]. These algorithms can be broadly classified into three categories: parametric, semi-parametric, and non-parametric approaches, based on their representational capacity and how it scales with the number of tasks and data samples. Table I shows a computational taxonomy of lifelong learners as a function of the sample size $n$ and the total task $T$, as well as the common scenario where the sample size is fixed per task and therefore proportional to the number of tasks, $n\propto T$. For time complexity, we only consider the training time complexity. The space complexity of the learner refers to the amount of memory space needed to store the learner [47]. We also study the representation capacity of these algorithms. Capacity is defined as the size of the subset of hypotheses that is achievable by the learning algorithm [48]. For a detailed discussion and assumptions behind the complexity analysis, see Appendix A-C.
IV Representation ensembling algorithms